# This work has been submitted to the IEEE Access for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible. xx.xxxx/ACCESS.2024.xxxxxxx

### v2

@@ Abstract
To adapt to real-world dynamics, intelligent systems need to assimilate new knowledge without catastrophic forgetting, where learning new tasks leads to a degradation in performance on old tasks. To address this, continual learning concept is proposed for enabling autonomous systems to acquire new knowledge and dynamically adapt to changing environments. Specifically, energy-efficient continual learning is needed to ensure the functionality of autonomous systems under tight compute and memory resource budgets (i.e., so-called autonomous embedded systems). Neuromorphic computing, with brain-inspired Spiking Neural Networks (SNNs), offers inherent advantages for enabling low-power/energy continual learning in autonomous embedded systems. In this paper, we comprehensively discuss the foundations and methods for enabling continual learning in neural networks, then analyze the state-of-the-art works considering SNNs. Afterward, comparative analyses of existing methods are conducted while considering crucial design factors, such as network complexity, memory, latency, and power/energy efficiency. We also explore the practical applications that can benefit from SNN-based continual learning and open challenges in real-world scenarios. In this manner, our survey provides valuable insights into the recent advancements of SNN-based continual learning for real-world application use-cases.
@@
Brain-inspired neuromorphic computing, spiking neural networks, event-based processing, continual learning, autonomous embedded systems.
@@ I Introduction
Humans continuously learn from their interactions with the environment throughout their lifetime [1], exhibiting remarkable cognitive flexibility of brains to enable simultaneous learning and retention of multiple skills. For instance, a person fluent in English can also master Spanish and French, each with unique grammatical structures and vocabulary, without losing proficiency in their native language. This ability is rooted in the brains’ delicate balance between plasticity and stability [2, 3]. Plasticity allows for the acquisition of new skills, while stability ensures that existing knowledge is preserved. Naturally, we expect Artificial Intelligence (AI) systems [4, 5, 6], wielding Neural Network (NN) algorithms [7], to develop a similar learning capability to effectively operate and adapt in real-world scenarios. While state-of-the-art AI systems excel in single task-based static environments, such as image recognition [8], face identification [9], and speech recognition [10], they often struggle in dynamic environments where data or tasks may change in structure, distribution, or characteristics over time. For instance, when a new task is encountered after a resource-intensive training, NNs are often retrained from scratch despite the high computational costs [11]. Therefore, robustness to multiple tasks and sequential experiences, remains a significant research challenge for AI systems. When faced with incremental learning of different tasks, most NNs underperform due to suffering from rapid performance degradation, a phenomenon known as Catastrophic Forgetting (CF) or interference [12, 13, 14].
@@
Numerous CL methods have been proposed to address CF in the conventional Artificial/Deep Neural Network (ANN/DNN) domain, which showed notable performance improvements [21, 22, 23, 24, 25, 26, 27]. However, most methods are resource-intensive, demanding additional memory and computational power. Therefore, DNN-based CL methods often fail to account for storage usage during incremental training, necessitating a significant addition of memory for parameters [28]. These conditions are not suitable for autonomous embedded systems, that use portable-battery and need quick adaptation to new data within limited resources. However, designing resource- and energy-efficient systems with CL capabilities (i.e., CL systems) is still a major challenge.
@@
Recently, the brain-inspired Neuromorphic Computing (NC) paradigm [29, 30, 31, 32, 33] has emerged as a promising field for enabling efficient and low-power information processing and decision-making, i.e., by emulating the architecture and functionality of brains with Spiking Neural Network (SNN) algorithms [34, 35, 36, 37, 38, 39]. The event-driven nature of SNNs supports energy-efficient learning of both static and non-static data streams [34, 35, 36, 37, 38, 39]. Furthermore, SNNs can perform unsupervised learning due to their bio-plausible learning rules (e.g., Spike-Timing-Dependent Plasticity (STDP) [40]), thus enabling continuous adaptation to dynamically changing environments and efficiently learning spatio-temporal data online without labels [41, 42]. In this manner, NC aligns seamlessly with the desired characteristics (desiderata) of CL, called Neuromorphic Continual Learning (NCL).
I-A Surveys on Continual Learning (CL)
To fully understand the foundations and recent advancements in CL, a comprehensive survey is imperative. Most of the existing CL surveys [16, 17, 43, 44, 45, 46, 47, 48, 49, 50, 51, 11, 52] have particularly focused on addressing CF in DNNs. Studies of [16] provided an up-to-date CL survey in DNNs. Other surveys focused on categorizing CL methods [43, 47, 44, 49], studying the challenges of forgetting in DNNs [11, 45, 50], reviewing computational requirements of CL (e.g., regularization) [11], discussing the benefits of forgetting in CL [45], providing experimental comparisons of 11 CL methods [50], discussing Online Continual Learning (OCL) paradigm [51], and discussing the importance of adopting brain-inspired data representations [53]. Although all these surveys provide insightful theoretical foundations and information regarding the recent advancements in CL for DNNs, most of them have not addressed CL with bio-plausible SNNs.
The remarkable CL capability of humans has sparked interest in the emerging of NCL paradigm, i.e., investigation on how SNN architectures and learning rules can be utilized to develop efficient CL systems. However, notable surveys on NCL still focused on exploring biological mechanisms for CL [34], and providing the background theories [54]. Therefore, they have not provided a holistic survey that provides a comprehensive view of NCL, ranging from theories to applications. Toward this, in this paper, we provide an in-depth review on preliminaries of CL and NC, state-of-the-art advancements of CL research in ANNs/DNNs and SNNs, and real-world applications for NCL (Fig. 2). Moreover, we also discuss the open research challenges, hence offering an updated perspective for future research directions in developing energy-efficient and adaptive AI systems. A comparison summary of our survey and the existing ones is presented in Table I.
Figure 2: Overview of our survey: CL in ANNs/DNNs vs. SNNs (i.e., NCL). DNNs rely on global learning, supervised training, and computationally intensive processing on conventional hardware, thus making them less suitable for OCL. In contrast, SNNs leverage event-based processing, local learning rules, and low-power neuromorphic hardware which enable energy-efficient CL, thus making them suitable for OCL.
@@ I-A Surveys on Continual Learning (CL)
TABLE I: A qualitative comparison of our survey and the existing ones. Note: “✓”, “+”, and “-” mean full, partial, and no consideration, respectively.
Work Year Research Objective Taxonomy Hardware Applications
Questions CL NCL Platforms
[17] 2019 - - - - -
[44] 2021 - - - +
[50] 2021 - - - -
[46] 2022 - - + - -
[47] 2022 - - + - +
[48] 2022 - - + - -
[49] 2022 - - + - +
[52] 2022 - - + - + -
[51] 2022 - - -
[43] 2023 - - - -
[45] 2023 - - - -
[34] 2023 - - - +
[53] 2023 - + - -
[11] 2024 - - + - -
[16] 2024 - - -
[54] 2024 + - -
Our Survey 2024
I-B Our Novel Contributions
Our survey addresses the key questions in developing CL with neuromorphic computing (SNNs), as the following.
  • What are the desiderata, evaluation scenarios, and performance metrics for CL?
  • What are the existing CL methods in the ML field to overcome CF, and how can they be systematically categorized?
  • What are the limitations of the existing CL methods?
  • Why are the energy-efficient CL algorithms crucial for resource-constrained autonomous embedded systems?
@@ I-B Our Novel Contributions
  • What techniques do the state-of-the-art propose to enable energy-efficient CL with SNNs (i.e., NCL)?
  • What are the real-world application use-cases that benefit from NCL?
  • What are the open challenges to be considered for enabling energy-efficient NCL?
@@
By addressing the above key questions, our paper makes the following major novel contributions:
@@
    We introduce the fundamentals of CL problems by explaining its basic formulation, desiderata, and possible learning scenarios. Furthermore, we highlight the OCL paradigm and its importance for practical applications.
@@
    We discuss the existing methods for enabling CL, which are pre-dominantly from DNN domain. Afterward, we provide different possible CL settings, and provides metrics for evaluating CL.
@@
    We discuss the background of NC (SNNs), then provide an up-to-date comprehensive review of state-of-the-art works for NCL.
  4. 4.
    We survey a range of real-world applications that benefit from the energy-efficient NCL. Toward the end, we discuss the open research challenges to provide a perspective on future research directions.

### v3

# This work has been accepted for publication in IEEE Access. xx.xxxx/ACCESS.2025.xxxxxxx

@@ Abstract
The challenging deployment of compute- and memory-intensive methods from Deep Neural Network (DNN)-based Continual Learning (CL), underscores the critical need for a paradigm shift towards more efficient approaches. Neuromorphic Continual Learning (NCL) appears as an emerging solution, by leveraging the principles of Spiking Neural Networks (SNNs) and their inherent advantages (e.g., sparse spike-driven operations and bio-plausible learning rules) for improving energy efficiency and performance, thereby enabling efficient CL algorithms (e.g., unsupervised learning approach) executed in dynamically-changed environments with resource-constrained computing systems. Though in its early stages, NCL is already a major research field with an increasing interest in novel SNN-based techniques for different CL methods (e.g., regularization-, replay-, and architecture-based). Motivated by the need for a holistic study of NCL, in this survey, we first provide a detailed background on CL, encompassing the desiderata, settings, metrics, scenario taxonomy, Online Continual Learning (OCL) paradigm, recent DNN-based methods proposed in the literature to address catastrophic forgetting (CF). Then, we analyze these methods based on their achieved CL desiderata, computational and memory costs, as well as network complexity, hence emphasizing the need for energy-efficient CL. After introducing the CL background and the energy efficiency challenges, we provide an extensive background of low-power neuromorphic computing systems including encoding techniques, neuronal dynamics, network architectures, learning rules, neuromorphic hardware processors, software and hardware frameworks, neuromorphic datasets, benchmarks, and evaluation metrics. Then, this survey comprehensively reviews and analyzes state-of-the-art works in the NCL field. The key ideas, implementation frameworks, and performance assessments (including CL, OCL, neuromorphic hardware compatibility aspects) are provided. This survey also covers several hybrid approaches that combine supervised and unsupervised learning paradigms and categorizes them into three main classes. It also covers optimization techniques including SNN operations reduction, weight quantization, and knowledge distillation. Then, this survey covers the progress of real-world NCL applications categorized into adaptive robots and autonomous vehicles with a wide range of use-cases i.e., object recognition, robotic arm control, cars and road lane detection, Simultaneous Localization and Mapping (SLAM), people detection and robotic navigation and provides their specific case-studies with empirical results. Finally, this paper provides a future perspective on the open research challenges for NCL, since the purpose of this study is to be useful for the wider neuromorphic AI research community and to inspire future research in bio-plausible OCL.
@@
Continual Learning (CL), Neuromorphic Computing, Spiking Neural Networks (SNNs), Neuromorphic Continual Learning (NCL), Event-based Processing, Energy Efficiency, Online Continual Learning (OCL), Catastrophic Forgetting (CF), Deep Neural Networks (DNNs), Artificial Intelligence (AI), Embedded AI Systems.
@@ I Introduction
The inherent cognitive flexibility of the human brain allows continuous learning from environmental interactions throughout life, exhibiting a remarkable ability to simultaneously acquire and retain multiple skills [1]. For instance, a person fluent in English can also master Spanish and French, each with unique grammatical structures and vocabulary, without losing proficiency in their native language. This ability is rooted in the brains’ delicate balance between plasticity and stability [2, 3]. Plasticity allows for the acquisition of new skills, while stability ensures that existing knowledge is preserved. Naturally, we expect Artificial Intelligence (AI) systems [4, 5, 6], wielding Neural Network (NN) algorithms [7], to develop a similar learning capability to effectively operate and adapt in real-world scenarios. While state-of-the-art AI systems excel in single task-based static environments, such as image recognition [8], face identification [9], and speech recognition [10], they often struggle in dynamic environments where data or tasks may change in structure, distribution, or characteristics over time. For instance, when a new task is encountered after a resource-intensive training, NNs are often retrained from scratch despite the high computational costs [11]. Therefore, robustness to multiple tasks and sequential experiences, remains a significant research challenge for AI systems. When faced with incremental learning of different tasks, most NNs underperform due to suffering from rapid performance degradation, a phenomenon known as Catastrophic Forgetting (CF) or interference [12, 13, 14].
@@
Numerous CL methods have been proposed to address CF in the conventional Artificial/Deep Neural Network (ANN/DNN) domain, which showed notable performance improvements [21, 22, 23, 24, 25, 26, 27]. However, most methods are resource-intensive, demanding additional memory and computational power. Therefore, ANN-based CL methods often fail to account for storage usage during incremental training, necessitating a significant addition of memory for parameters [28]. These conditions are not suitable for embedded AI systems, that use portable-battery and need quick adaptation to new data within limited resources. However, designing resource- and energy-efficient systems with CL capabilities (i.e., CL systems) is still a major challenge.
@@
Recently, the brain-inspired Neuromorphic Computing (NC) paradigm [29, 30, 31, 32, 33, 34] has emerged as a promising field for enabling efficient and low-power information processing and decision-making, i.e., by emulating the architecture and functionality of brains with Spiking Neural Network (SNN) algorithms [35, 36, 37, 38, 39, 40]. The event-driven nature of SNNs supports energy-efficient learning of both static and non-static data streams [35, 36, 37, 38, 39, 40]. Furthermore, SNNs can perform unsupervised learning due to their bio-plausible learning rules (e.g., Spike-Timing-Dependent Plasticity (STDP) [41]), thus enabling continuous adaptation to dynamically changing environments and efficiently learning spatio-temporal data online without labels [42, 43]. In this manner, NC aligns seamlessly with the desired characteristics (desiderata) of CL, called Neuromorphic Continual Learning (NCL).
@@
TABLE I: Qualitative comparison of our survey and the existing ones. Note: ”✓”, $\approx$”, and ”$\times$” mean full, partial, and no consideration, respectively.
Surveys
Year
Pages #
References #
Scope
Taxonomy
HW Platforms
SW Frameworks
Architectures
Benchmarks
Datasets
Metrics
Challenges
Quant. Analysis
Case Studies
CL NCL
[17] 2019 $21^{*}$ 207 $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
[44] 2021 25 344 $\times$ $\times$ $\times$ $\times$ $\times$
[45] 2021 18 102 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[46] 2022 18 159 $\times$ $\times$ $\times$ $\times$ $\times$ $\approx$ $\approx$
[47] 2022 18 124 $\times$ $\times$ $\times$ $\times$ $\times$
[48] 2022 $32^{*}$ 249 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[49] 2022 7 113 $\times$ $\approx$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[50] 2022 6 78 $\times$ SNNs $\times$ $\approx$ $\times$ $\approx$ $\times$ $\times$ $\times$
[51] 2022 30 79 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[52] 2023 15 281 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\approx$
[35] 2023 11 83 $\times$ $\times$ $\times$ $\approx$ $\approx$ $\times$ $\times$
[53] 2023 18 173 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[54] 2024 16 289 $\times$ $\times$ $\times$ $\times$ $\times$
[11] 2024 13 177 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[16] 2024 20 527 $\times$ $\times$ $\times$ $\times$ $\times$ $\approx$ $\times$ $\times$
[55] 2024 $28^{*}$ 66 $\approx$ $\times$ $\times$ $\times$ $\times$ $\approx$ $\times$ $\times$
[56] 2024 $25^{*}$ 304 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[57] 2024 7 80 $\times$ $\times$ $\times$ $\times$ $\approx$ $\times$
[58] 2024 17 167 $\times$ $\times$ $\approx$ $\approx$ $\approx$ $\approx$ $\times$ $\times$
[59] 2024 14 144 $\times$ $\times$ $\times$ $\times$ $\times$
[60] 2025 27 233 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
This work 2025 39 413
  • * Single-column pages.
I-A Surveys on CL and NCL
To fully understand the foundations and recent advancements in CL and NCL, a comprehensive survey is imperative, unlike other surveys that focus on a single area (either CL, NCL, bio-plausible learning rules, or applications). Table I compares surveys studying multiple aspects of CL, NCL or SNNs, like our work. Most of the existing CL surveys [16, 17, 44, 52, 46, 54, 47, 48, 49, 45, 51, 11, 50] have particularly focused on addressing CF in DNNs. For instance, studies in [16] provided an up-to-date CL survey in DNNs. Meanwhile, other surveys focused on categorizing CL methods [54, 47, 44, 49], studying the challenges of forgetting in DNNs [11, 52, 45], reviewing computational requirements of CL (e.g., regularization) [11], discussing the benefits of forgetting in CL [52], providing experimental comparisons of 11 CL methods [45], discussing Online Continual Learning (OCL) paradigm [51], and discussing the importance of adopting brain-inspired data representations [53]. A more recent line of surveys have also explored CL in large pre-trained models, focusing on paradigms such as prompt tuning, adapters, and distillation for mitigating forgetting in large language models (LLMs) [56, 57, 58, 59, 60]. While these methods offer solutions for foundation models, they are fundamentally different from resource-constrained, biologically-inspired approaches like NCL for enabling energy-efficient CL and OCL, which remain underexplored in existing surveys. In particular, our survey differs from the study [16] in the following ways (see overview in Table I and an in-depth comparison in Table II).
  • Focus: Survey [16] focused on categorizing DNN-based CL methods, and provided theoretical concepts of stability-plasticity trade-offs, generalizability in CL, and application domains. Our survey builds on that foundation to highlight the limitations of DNN-based CL methods, i.e., computational and memory overhead for resource-constrained systems. Motivated by these challenges, it provides a specialized perspective on state-of-the-art NCL methods, categorizing the existing literature into: 1) enhancements on unsupervised STDP learning, 2) predictive coding, 3) active dendrites, 4) bayesian continual learning, 5) architecture, 6) replay, 7) regularization and 8) hebbian learning based methods, along with optimization techniques proposed by these works, and SNN-specific application use-cases for enabling energy-efficient CL and OCL in embedded AI systems.
  • Comparative Quantitative Analysis: Our survey distinguishes itself from [16] by providing a comparative quantitative analysis considering design factors (i.e., network complexity) and evaluation metrics (i.e., accuracy, memory footprint, latency, power/energy usage) of the reviewed NCL methods with relevant DNN-based CL methods. It also provides a detailed quantitative trade-off analysis (accuracy vs. efficiency) for various NCL methods. Additionally, it provides a comparison of prominent neuromorphic datasets and a comparative analysis of NCL frameworks with standard DNN-based CL frameworks, which are not provided in [16].
  • Application Domains: Our survey systematically analyzes the current progress of NCL in real-world applications by categorizing existing works into: 1) adaptive robots, and 2) autonomous vehicles/agents. It provides specific case studies with empirical results (latency, power/energy, accuracy), and OCL capability assessment, to demonstrate the real-world feasibility of NCL in embedded AI systems. Furthermore, it reports the datasets, learning rules, software/hardware platforms, and discusses hardware deployment challenges. These aspects are not covered by [16] survey, as it mainly focuses on the scenario complexity and task-specific challenges.
TABLE II: Comparison of our survey with study [16]. Note: ”✓”, $\approx$”, and ”$\times$” mean full, partial, and no consideration, respectively.
Aspect This Work [16]
Taxonomy & categorization NCL methods & their real-world use-cases CL methods
Neuromorphic datasets & SNN-based evaluation metrics $\times$
SNN-based CL, OCL capability, & hardware compatibility $\times$
Hardware deployment constraints $\approx$
Quantitative trade-off analysis $\times$
Comparative analysis of NCL frameworks with standard DNN-based CL frameworks $\times$
Comparative quantitative analysis of NCL methods with DNN-based CL methods $\times$
The remarkable CL capability of humans has sparked interest in the emerging of NCL paradigm, i.e., investigation on how SNN architectures and learning rules can be utilized to develop efficient CL systems. For instance, studies in [35] explored biological mechanisms for NCL and emphasized the importance of CL, but did not provide a comparative quantitative analysis of the reviewed NCL methods with relevant DNN-based CL methods. It also did not provide details of neuromorphic datasets or a taxonomy that categorizes the recent literature on NCL methods and SNN-specific real-world application use-cases. The other survey [55] provided the background theories on sparse and predictive coding and their relation to Hebbian and bio-plausible learning, hence it only focused on learning rules and application aspect of NCL. It did not cover the aspects such as OCL capability, software and hardware frameworks, hardware compatibility, neuromorphic datasets & benchmarks, performance and computational efficiency trade-offs of existing NCL works. Therefore, a unified survey that covers the foundations of CL, recent advances in both the DNN and SNN domains, provides a comparative and application-driven view of NCL with detailed NCL design and optimization aspects, assessing suitability for OCL and neuromorphic hardware compatibility, has not been provided.
Figure 2: Overview of our survey: CL in ANNs/DNNs and SNNs (i.e., NCL). ANNs/DNNs rely on global learning, supervised training, and computationally intensive processing on conventional hardware, thus making them less suitable for OCL. SNNs leverage event-based processing, local learning rules, and low-power neuromorphic hardware which enable energy-efficient CL, thus making them suitable for OCL.
I-B Scope and Contribution
Our survey addresses the following key questions.
  • Q1:
    What are the CL desiderata, settings, evaluation metrics, and taxonomy of scenarios?
  • Q2:
    What are the different categories of existing CL methods in the ANN/DNN domain, how they overcome CF, and their achieved desiderata?
  • Q3:
    What are the deployment challenges of these CL methods (i.e., computational/memory overhead) that necessitate energy-efficient approaches?
  • Q4:
    What is the technical background of low-power neuromorphic systems, current software and hardware frameworks/platforms, datasets, benchmarks and evaluation metrics?
  • Q5:
    How do neuromorphic frameworks compare with standard deep learning (DL) and CL frameworks?
  • Q6:
@@ I-B Scope and Contribution
    What methods do the state-of-the-art propose to enable energy-efficient CL with SNNs (i.e., NCL), their OCL capability, and which efficiency enhancement techniques they propose?
  • Q7:
    What is the compatibility of reviewed methods with neuromorphic hardware platforms, and their on-chip learning capabilities?
  • Q8:
    Is there a standardized evaluation framework or metrics for benchmarking NCL and what additional metrics can be used for evaluating NCL methods?
  • Q9:
    What are the hardware deployment challenges of NCL methods considering accuracy-efficiency trade-offs?
  • Q10:
    What are the specific use-cases in adaptive robots and autonomous vehicles that benefit from NCL, their OCL capability, and quantitative performance analysis?
  • Q11:
    What are the identified open challenges and proposed future research directions?
@@
By addressing the above key questions, our paper covers all the key aspects discussed above and makes the following contributions (see overview in Fig. 2):
@@
    We provide the preliminaries of CL including formulation, desiderata, settings, taxonomy of learning scenarios, evaluation metrics, and OCL paradigm. We discuss how different categories of existing CL methods which are pre-dominantly from ANN/DNN domain overcome CF, and which of the desiderata they achieve. Furthermore, we emphasize their challenges (computational/memory overhead) for practical applications in resource-constrained systems, outlining the need for energy-efficient approaches.
@@
    We discuss the technical background of low-power neuromorphic computing systems covering the key aspects (i.e., encoding schemes, neuronal dynamics, learning rules, software/hardware frameworks, neuromorphic processor architectures, an elaborate description of neuromorphic datasets, benchmarks & metrics), suggesting the need for standardized NCL benchmarks and additional metrics. We provide a comparative analysis of NCL frameworks with conventional DNN-based CL frameworks. Then, we provide a comprehensive review of state-of-the-art works for NCL, categorizing them into: 1) enhancements on unsupervised STDP learning, 2) predictive coding, 3) active dendrites, 4) bayesian continual learning, 5) architecture, 6) replay, 7) regularization and 8) hebbian learning based methods, assessing their CL performance, OCL capabilities, and neuromorphic hardware compatibility. We discuss hybrid approaches that combine supervised and unsupervised learning paradigms to address CF and improve CL performance. We categorize these approaches into three main classes: 1) self-supervised pre-training hybrids, 2) STDP + supervised learning hybrids, and 3) generative-discriminative hybrid models. We discuss efficiency enhancement techniques for NCL proposed by state-of-the-art including SNN operations reduction, quantization, and knowledge distillation. Moreover, we discuss the performance and computational efficiency trade-offs of NCL considering hardware implementation constraints and provide a comparative quantitative analysis considering design factors and evaluation metrics of the reviewed NCL with relevant DNN-based CL methods.
@@
    We categorize the current progress of NCL in real-world applications into: 1) adaptive robots, and 2) autonomous vehicles/agents, studying a range of emerging use-cases including object recognition, robotic arm control, cars and road lane detection, Simultaneous Localization and Mapping (SLAM), people detection and robotic navigation, and report their specific case-studies with empirical results, implementation frameworks, and OCL capabilities to assess real-world feasibility. Toward the end, we discuss the identified open research challenges and provide a perspective on future research directions.
Table III summarizes the abbreviations used throughout the paper for clarity and consistency.
TABLE III: List of Abbreviations
Abbreviation Full Form
CL Continual Learning
DL Deep Learning
ANN Artificial Neural Network
DNN Deep Neural Network
CNN Convolutional Neural Network
NNs Neural Networks
CF Catastrophic Forgetting
NCL Neuromorphic Continual Learning
SNN Spiking Neural Network
CSNN Convolutional SNN
OCL Online Continual Learning
TA Task Agnostic
FWT Positive Forward Transfer
BWT Positive Backward Transfer
RL Reinforcement Learning
STDP Spike-Timing-Dependent Plasticity
SLAM Simultaneous Localization and Mapping
Task-IL Task-Incremental Learning
Domain-IL Domain-Incremental Learning
Class-IL Class-Incremental Learning
DTA Discrete Task Agnostic
CTA Continuous Task Agnostic
OWTA Open-World Task Agnostic
SSL Self-Supervised Learning
CPT Continual Pre-Training
BP Backpropagation
EWC Elastic Weight Consolidation
FIM Fisher Information Matrix
SI Synaptic Intelligence
GEM Gradient Episodic Memory
GPM Gradient Projection Memory
MAS Memory Aware Synapses
RWalk Riemannian Walk
MLP Multi-layer Perceptron
FC Fully-Connected
CFN Controlled Forgetting Network
ASP Adaptive Synaptic Plasticity
ViT Vision Transformer
KD Knowledge Distillation
ER Experience Replay
HW Hardware
SW Software
TTFS Time-to-first Spike
HH Hodgkin–Huxley
IF Integrate-and-Fire
LIF Leaky IF
AdExIF Adaptive Exponential IF
SDSP Spike-Driven Synaptic Plasticity
R-STDP Reward-Modulated STDP
PES Prescribed Error Sensitivity
BPTT Backpropagation Through Time
SG Surrogate Gradient
STBP Spatio-Temporal Backpropagation
DECOLLE Deep Continuous Local Learning
ST-LRA Spike-Triggered Local Representation Alignment
BrainCog Brain-inspired Cognitive Intelligence Engine
CPU Central Processing Unit
GPU Graphic Processing Unit
MCU Microcontroller Unit
TPU Tensor Processing Unit
CMOS Complementary Metal-Oxide-Semiconductor
FPGA Field-Programmable Gate Array
ASIC Application-Specific Integrated Circuit
ROLLS Reconfigurable On-line Learning Spiking
Neuromorphic Processor
PIM Processing-In-Memory
CIM Compute-In-Memory
NVM Non-Volatile Memory
RRAM Resistive Random Access Memory
MRAM Magnetic RAM
PCM Phase Change Memory
TABLE III: Continued
Abbreviation Full Form
DVS Dynamic Vision Sensor
ATIS Asynchronous Time-based Image Sensor
DAS Dynamic Audio Sensor
NVS Neuromorphic Vision Sensor
BAE Biologically Plausible Auditory Encoding
mAP Mean Average Precision
MAE Mean Absolute Error
$MAE_{L}$ MAE for Localisation
$MAE_{M}$ MAE for Mapping
MSE Mean Square Error
RMSE Root Mean Square Error
AOC Average of top-1 accuracy over classes
ITAE Integral of Time-weighted Absolute Error
SNR Signal-to-Noise Ratio
SpNCN Spiking Neural Coding Network
DSD-SNN Dynamic Structure Development of SNN
SOR-SNN Self-Organized Regulation SNN
LR Latent Representation/Replay
HLOP Hebbian Learning-based Orthogonal Projection
NEF Neural Engineering Framework
NSM Neural State Machine
IK Inverse Kinematics
PID Proportional Integral Derivative
PES Prescribed Error Sensitivity
WCE Weighted Binary Cross Entropy
IoU Intersection over Union
FMCW Frequency Modulated Continuous Wave
HNN Hybrid Neural Network (SNN-ANN)
FPS/W Frames Per Second per Watt
DoF Degrees of Freedom
FID Fréchet distance